# CoD: A Diffusion Foundation Model for Image Compression

### v1

@@ 2.1 Pipeline
Entropy bottleneck. It is designed to maintain an ultra-low bitrate for $y$, encouraging the diffusion model to learn strong generative capabilities. This can be implemented as scalar quantization with entropy models [hyperprior], vector quantization [vqgan], or finite scalar quantization [fsq]. For simplicity, we use vector quantization with a codebook size of $N=2^{4}=16$, corresponding to $4\mathrm{bpp}/(32\times 32)=0.0039\text{ bpp}$.
@@
In the latent space, since VAE latents are already at $1/8$ resolution, we apply a $2\times 2$ patch embedding to downsample them to $1/16$, and upsample the output of DDT head back to $1/8$ before passing to the VAE decoder. In the pixel space, we directly use a $16\times 16$ patch embedding to map the noised images to $1/16$ resolution, and adopt the pixel DDT head from [pixnerd], where each feature predicts a neural field that reconstructs a corresponding $16\times 16$ patch. Since the DiT backbone runs at $1/16$ resolution of original images, pixel-space CoD has a similar computation complexity as in latent space. Details are in the supplementary material.
@@ 2.3 Implementation Details
Loss Function. For the information bottleneck, we adopt the codebook commitment loss $\mathcal{L}_{\mathrm{C}}$ [vqgan]. Diffusion training employs the rectified flow loss $\mathcal{L}_{\mathrm{RF}}$ and the representation alignment loss $\mathcal{L}_{\mathrm{REPA}}$ with DINOv2 [dino_v2]. Additionally, an auxiliary head (Figure 2) reconstructs both original pixels and DINOv2 features from the condition $c$ to enhance condition learning. Further analysis of this auxiliary loss $\mathcal{L}_{\mathrm{aux}}$ is in the supplementary material. The overall objective is
@@ 3 Analysis on Foundation Model
In this section, we conduct an in-depth analysis of CoD to investigate its performance and underlying characteristics. More analysis are in the supplementary material.
@@ 4 Downstream Compression Performance
In this section, we conduct a comprehensive evaluation of CoD when integrated into downstream compression frameworks. Since most diffusion-based codecs do not release their training code, we employ the zero-shot compression method DiffC [diffc] on CoD for comparison.
Benchmarks and Baselines. Experiments are conducted on the Kodak [kodak], CLIC2020 test set [clic] and Div2K [div2k]. All images are resized and center-cropped to a resolution of $512\times 512$. We report distortion using PSNR and perceptual quality using LPIPS [lpips], DISTS [dists], and FID [fid]. Following [ddcm], FID is computed on extracted patches ($64\times 64$ for Kodak and $128\times 128$ for CLIC2020). Bitrate is measured in bits per pixel (bpp). In addition to diffusion codecs (DiffC, DDCM, PerCo (SD), and OSCAR; see Section 3.1), we further compare against the latent-space codec DiffEIC [diffeic] and pixel-space codecs including the traditional codec VTM [vvc] and perceptual codecs HiFiC [hific], MS-ILLM [illm], CDC [cdc] and TACO [taco]. We also discuss StableCodec [stablecodec] and OneDC [onedc] which are built on pretrained one-step diffusion models in the supplementary.
@@ 6 Conclusion and Limitation
This document provides the supplementary material to Compression-oriented Diffusion foundation model, CoD.
@@ Appendix A Towards 64-Bit Image Compression
Figure 11: Rate-Distortion comparison for 64-bit CoD and Stable Diffusion based codecs on Kodak at $512\times 512$.
@@ Appendix B Towards One-Step Diffusion Compression
Preliminary. Recently, one-step diffusion models have demonstrated effectiveness in generative image compression. Unlike multi-step diffusion, which optimizes each diffusion timestep separately, one-step diffusion can be trained end-to-end, enabling better overall performance. OSCAR [oscar] fine-tuned a multi-step Stable Diffusion to function as a one-step image codec. Subsequently, one-step diffusion foundation models have been explored for generative image compression. StableCodec [stablecodec] further leverages a one-step SD-Turbo as the foundation model, achieving significantly improved performance. Similarly, OneDC [onedc] employs a one-step DMD-distilled [dmd] Stable Diffusion as the foundation model for one-step diffusion compression. The success of these approaches motivates our exploration of CoD in the context of one-step diffusion foundation models.
@@
Distilling a CoD. Following OneDC, we train a one-step CoD using a combination of pixel-space loss and DMD loss:
@@
$\mathcal{L_{\mathrm{OS}}}=\mathcal{L}_{1}+\mathcal{L}_{\mathrm{LPIPS}}+\lambda\,\cdot\,\mathcal{L}_{\mathrm{REPA}}+\zeta\,\cdot\,\mathcal{L}_{\mathrm{DMD}},$ (10)
where $\mathcal{L}_{1}$ and $\mathcal{L}_{\mathrm{LPIPS}}$ are the pixel-space L1 loss and the VGG [lpips] LPIPS loss, and $\mathcal{L}_{\mathrm{REPA}}$ is the representation alignment loss. We set $\lambda=0.5$ and $\zeta=0.625$. The condition encoder and decoder are kept fixed while tuning the diffusion module, as optimizing the condition during training could introduce discrepancies between the teacher and student models, potentially causing catastrophic training failure.
@@
For simplicity, we optimize only the pixel-space CoD, allowing pixel-space losses to be computed directly. Following the DMD procedure, we update the one-step CoD every 10 steps, while the remaining 9 steps train the fake score diffusion model. The model is trained for 50K steps (i.e., 5K steps for the one-step CoD) with a batch size of 32, which takes approximately 12 hours on 4 A100 GPUs (48 A100 GPU hours in total).
Figure 12: Evaluation of one-step CoD on Kodak at $512\times 512$.
Evaluation. We evaluate our one-step CoD on Kodak in Figure 12. With very limited training steps, one-step CoD achieves impressive reconstruction quality. Although its detail is not as clear as multi-step CoD, its hue is more similar to the original image, and some structures exhibit higher fidelity. Compared to the DMD-distilled diffusion codec OneDC, one-step CoD has better perceptual quality. The rate-distortion curves tested on Kodak are shown in Figure 14, where one-step CoD showcases better PSNR and LPIPS values than multi-step CoD, while DISTS and FID are slightly worse. Compared to other one-step diffusion codecs, one-step CoD has comparable or better perceptual metrics and a higher PSNR.
Note. In this paper, we train one-step CoD with limited steps (5K generator optimization steps), mainly to explore its properties rather than delivering a fully optimized one-step diffusion codec. Notably, training a one-step Stable Diffusion using DMD costs approximately 1664 A100 GPU hours, whereas we use 48 A100 hours. We believe longer training, a larger batch size, or larger-scale datasets can further boost its performance.
@@ Appendix C Towards Perceptual Supervision via CoD
Methods PSNR LPIPS* DISTS FID
@@
+ CoD-based DMD loss 21.10 dB 0.402 0.255 80.1
    LPIPS is measured with AlexNet. MS-ILLM is trained with LPIPS (Alex), while we follow a more common practice LPIPS (VGG) to avoid train-test overfitting, explaining the minor LPIPS gain.
@@
Evaluation. MS-ILLM is typically pretrained using the MSE loss, after which the encoder and entropy model are fixed while the decoder is finetuned with perceptual losses such as LPIPS and GAN [gan] loss. Following this paradigm, we finetune only the decoder of MS-ILLM, replacing the GAN loss with the CoD-based DMD loss. Remarkably, after optimizing the decoder for just 2K steps with a batch size of 32 (20K steps total including fake score learning, requiring only about 14 A100 GPU hours), MS-ILLM demonstrates substantially improved realism, as shown in Table 1. The CoD-based DMD loss enhances perceptual metrics, notably reducing FID from 92.5 to 80.1. Several visual examples are presented in Figure 13, where the CoD-based DMD loss markedly reduces artifacts and yields clearer edges and finer details. Since the encoder is fixed, the encoded information remains heavily distorted under MSE-only optimization, resulting in poor overall reconstruction. We expect that jointly tuning the entire network with the CoD-based DMD loss could further improve perceptual quality.
@@
Figure 13: Finetuning MS-ILLM decoder with CoD-based DMD loss.
Figure 14: Rate-Distortion comparison for one-step CoD and other one-step diffusion codecs on Kodak at $512\times 512$.
Figure 15: More visualization results on CLIC 2020 test set [clic].
Appendix D Exploring $\mathcal{X}$-Prediction Pixel Diffusion
@@
Pixel-space CoD with JiT. CoD does not require any specialized network design for the diffusion model, so we can directly replace our pixel-space diffusion model with JiT. We refer to this variant as CoD (pixel-$\mathcal{X}$) to distinguish it from our velocity-based pixel-space CoD. Following JiT, CoD (pixel-$\mathcal{X}$) predicts denoised pixels while being trained with a velocity loss, maintaining consistency with our original training objective.
Comparison. We train both CoD (pixel) and CoD (pixel-$\mathcal{X}$) at $256\times 256$ resolution under identical conditions for 200K steps and evaluate them on Kodak. As shown in Table 2, compared to $\mathcal{V}$-prediction, we note $\mathcal{X}$-prediction improves PSNR, LPIPS and DISTS by a large margin, while achieving a comparable FID score. It demonstrates the superiority of $\mathcal{X}$-prediction in terms of reconstruction fidelity. In future, we will explore $\mathcal{X}$-prediction for more advanced pixel-space diffusion codecs.
Method @ 0.0156 bpp PSNR LPIPS DISTS FID
CoD (pixel) 15.00 dB 0.378 0.224 92.7
CoD (pixel-$\mathcal{X}$) 16.99 dB 0.348 0.207 93.4
Table 2: Ablation study for $\mathcal{X}$-prediction on Kodak at $256\times 256$. Each model is trained for 200k steps for comparison.
Appendix E Additional Details and Results

### v2

@@ 2.1 Pipeline
Entropy bottleneck. It is designed to maintain an ultra-low bitrate for $y$, encouraging the diffusion model to learn strong generative capabilities. This can be implemented as scalar quantization with entropy models [hyperprior], vector quantization [vqgan], or finite scalar quantization [fsq]. For simplicity, we use vector quantization with a codebook size of $N=2^{4}=16$, corresponding to $4\text{ bits}/(32\times 32)=0.0039\text{ bpp}$.
@@
In the latent space, since VAE latents are already at $1/8$ resolution, we apply a $2\times 2$ patch embedding to downsample them to $1/16$, and upsample the output of DDT head back to $1/8$ before passing to the VAE decoder. In the pixel space, we directly use a $16\times 16$ patch embedding to map the noised images to $1/16$ resolution, and adopt the pixel DDT head from [pixnerd], where each feature predicts a neural field that reconstructs a corresponding $16\times 16$ patch. Since the DiT backbone runs at $1/16$ resolution of original images, pixel-space CoD has a similar computation complexity as in latent space. Details are in the appendix.
@@ 2.3 Implementation Details
Loss Function. For the information bottleneck, we adopt the codebook commitment loss $\mathcal{L}_{\mathrm{C}}$ [vqgan]. Diffusion training employs the rectified flow loss $\mathcal{L}_{\mathrm{RF}}$ and the representation alignment loss $\mathcal{L}_{\mathrm{REPA}}$ with DINOv2 [dino_v2]. Additionally, an auxiliary head (Figure 2) reconstructs both original pixels and DINOv2 features from the condition $c$ to enhance condition learning. Further analysis of this auxiliary loss $\mathcal{L}_{\mathrm{aux}}$ is in the appendix. The overall objective is
@@ 3 Analysis on Foundation Model
In this section, we conduct an in-depth analysis of CoD to investigate its performance and underlying characteristics. More analysis are in the appendix.
@@ 4 Downstream Compression Performance
In this section, we conduct a comprehensive evaluation of CoD when integrated into downstream compression frameworks. Since most diffusion-based codecs do not release their training code, we employ the zero-shot compression method DiffC [diffc] on CoD for comparison. Additional downstream results are presented in the appendix, e.g., finetuning one-step CoD achieves real-time and SOTA performance (Section B), and CoD-based perceptual loss that significantly enhances MS-ILLM [illm] performance (Section C).
Benchmarks and Baselines. Experiments are conducted on the Kodak [kodak], CLIC2020 test set [clic] and Div2K [div2k]. All images are resized and center-cropped to a resolution of $512\times 512$. We report distortion using PSNR and perceptual quality using LPIPS [lpips], DISTS [dists], and FID [fid]. Following [ddcm], FID is computed on extracted patches ($64\times 64$ for Kodak and $128\times 128$ for CLIC2020). Bitrate is measured in bits per pixel (bpp). In addition to diffusion codecs (DiffC, DDCM, PerCo (SD), and OSCAR; see Section 3.1), we further compare against the latent-space codec DiffEIC [diffeic] and pixel-space codecs including the traditional codec VTM [vvc] and perceptual codecs HiFiC [hific], MS-ILLM [illm], CDC [cdc] and TACO [taco]. We also discuss StableCodec [stablecodec] and OneDC [onedc] which are built on pretrained one-step diffusion models in the appendix.
@@ 6 Conclusion and Limitation
Change Log
  • v1 (2025-11-24): Initial submission.
  • v2 (2025-12-10): Updated algorithms and results for finetuned one-step diffusion codec (App. B), CoD-based perceptual supervision (App. C), analysis on different downstream codecs (App. D), and $\mathcal{X}$-prediction (App. E).
@@
This document provides the supplementary material to Compression-oriented Diffusion foundation model, CoD. Beyond additional training details and extended results, we further investigate its characteristics and explore additional downstream codec applications.
The key conclusions are summarized below:
  • Extreme 64-bit compression: CoD compresses image into only 64 bits while preserving correct semantics.
  • One-step diffusion codec: Finetuning one-step CoD achieves real-time and SOTA performance, comparable to StableCodec [stablecodec] and OneDC [onedc].
  • CoD as perceptual loss: CoD-based perceptual loss significantly improves MS-ILLM [illm].
  • $\mathcal{X}$-prediction advantage: Using $\mathcal{X}$-prediction yields better performance than $\mathcal{V}$-prediction for CoD (pixel).
@@ Appendix A Towards 64-Bit Image Compression
Figure 11: Rate-Distortion curves for 64-bit CoD and Stable Diffusion based codecs on Kodak at $512\times 512$.
Figure 12: Rate-Distortion curves for finetuned one-step CoD and other one-step diffusion codecs on Kodak at $512\times 512$.
Figure 13: Evaluation of one-step CoD on Kodak at $512\times 512$.
Figure 14: Rate-Distortion curves for MS-ILLM [illm] and finetuning MS-ILLM decoder with CoD-based DMD loss on Kodak at $512\times 512$.
@@ Appendix B Towards One-Step Diffusion Compression
B.1 Preliminary
Recently, one-step diffusion models have demonstrated effectiveness in generative image compression. Unlike multi-step diffusion, which optimizes each diffusion timestep separately, one-step diffusion can be trained end-to-end, enabling better overall performance. OSCAR [oscar] fine-tuned a multi-step Stable Diffusion to function as a one-step image codec. StableCodec [stablecodec] leverages a one-step SD-Turbo as the foundation model, achieving significantly improved performance. Similarly, OneDC [onedc] employs a one-step DMD-distilled [dmd] Stable Diffusion as the foundation model for one-step diffusion compression. The success of these approaches motivates our exploration of CoD in the context of one-step diffusion foundation models.
@@ B.1 Preliminary
B.2 Distilling a CoD
We train a one-step CoD using a combination of losses and multi-step CoD-based DMD loss:
@@ B.2 Distilling a CoD
$\mathcal{L_{\mathrm{OS}}}=\mathcal{L}_{\mathrm{pixel}}+\mathcal{L}_{\mathrm{perc}}+\lambda\,\cdot\,\mathcal{L}_{\mathrm{REPA}}+\beta\,\cdot\,\mathcal{L}_{\mathrm{C}},$ (10)
where $L_{\mathrm{pixel}}=\mathcal{L}_{1}+\mathcal{L}_{\mathrm{LPIPS}}$ consists of pixel-space L1 loss and the LPIPS [lpips] loss, $L_{\mathrm{perc}}=2\,\cdot\,\mathcal{L}_{\mathrm{DMD}}+0.01\,\cdot\,\mathcal{L}_{\mathrm{GAN}}$ consists of a DMD loss and a patch GAN loss [vqgan], $\mathcal{L}_{\mathrm{REPA}}$ is the representation alignment loss and $\mathcal{L}_{\mathrm{C}}$ is th codebook commitment loss. We set $\lambda=0.5$ and $\beta=0.25$.
@@
For simplicity, we optimize only the pixel-space CoD, allowing pixel-space losses to be computed directly. Following the DMD procedure, we update the one-step CoD every 10 steps, while the remaining 9 steps train the fake score diffusion model. The model is trained for 100K steps (i.e., 10K steps for the one-step CoD) with a batch size of 32, which takes approximately 96 A100 GPU hours in total, which is much less than 1664 A100 GPU hours of DMD-based Stable Diffusion distillation. The learning rates for one-step CoD and fake score network are set to $10^{-5}$.
B.3 Finetuning to Higher Bitrates
In the previous section, we distilled the one-step CoD model at 0.0039 bpp. We further observe that the distilled model can be efficiently adapted to higher bitrates. Following PerCo and OSCAR, we adjust both the downsampling scale in the condition encoder and the codebook size according to the target bitrate. Specifically, we train two variants: 0.0312 bpp with $16\times$ downsampling and codebook size 256, and 0.125 bpp with $8\times$ downsampling and codebook size 256.
Stage I: initialization. We randomly set the parameters of the condition encoder, codebook, and condition decoder, while loading the pre-trained multi-step CoD weights for the diffusion module. During this warm-up phase, the diffusion module is trained using LoRA [lora] with rank 32. To accelerate convergence, we remove the $L_{\mathrm{perc}}$ term from $L_{\mathrm{OS}}$ and train for 200K steps with a learning rate of $10^{-4}$.
Stage II: Fine-tuning. We then fine-tune all model components jointly using the full $L_{\mathrm{OS}}$ for another 100K steps with a batch size of 32 and a learning rate of $10^{-5}$. With perceptual supervision restored, the fine-tuned one-step CoD achieves significantly improved performance at higher bitrates.
B.4 Evaluation
Comparison with one-step codecs. We evaluate the fine-tuned one-step CoD on the Kodak dataset (Figure 12). It yields competitive reconstruction quality compared with SOTA one-step diffusion codecs such as OneDC and StableCodec, particularly at ultra-low bitrates. Notably, our method does not rely on any pretrained components from Stable Diffusion (e.g., one-step models or SD-VAEs) to reach this performance. We currently adopt fixed-length coding for the codebook indices for simplicity. Further gains can be achieved by introducing an entropy model or applying latent compression [glc]. Visual comparisons in Figure 13 further show that one-step CoD preserves higher fidelity than both multi-step CoD and OneDC.
Complexity Analysis. DiT injects conditios and timesteps through AdaLN-Zero layers, which account for more than 200M parameters in CoD. However, CoD concatenates compression conditions directly with the noised input along the channel dimension instead of injecting them through AdaLN-Zero (see Section 2.1). Under the one-step setting, the timestep is fixed to 0, making all AdaLN-Zero outputs constant and therefore redundant. Thus, we precompute these constants and remove the AdaLN-Zero modules to further reduce computation and memory access. As shown in Table 5, one-step CoD processes a $512\times 512$ image in only 25.2 ms, which is even faster than a single step of Stable Diffusion used in OneDC and StableCodec. This demonstrates its potential in real-time application.
@@ Appendix C Towards Perceptual Supervision via CoD
Methods PSNR LPIPS DISTS FID
@@
+ CoD-based DMD loss 21.08 dB 0.376 0.248 80.92
@@
Figure 15: Finetuning MS-ILLM decoder with CoD-based DMD loss.
Figure 16: Rate-Distortion curves for different downstream codecs built on pixel-space CoD on Kodak at $512\times 512$.
Evaluation. MS-ILLM is typically pretrained using the MSE loss, after which the encoder and entropy model are fixed while the decoder is finetuned with perceptual losses such as LPIPS and GAN [gan] loss. Following this paradigm, we finetune only the decoder of MS-ILLM, replacing the GAN loss with the CoD-based DMD loss. Remarkably, after optimizing the decoder for just 2K steps with a batch size of 32 (20K steps total including fake score learning), MS-ILLM demonstrates substantially improved perceptual quality across all metrics, as shown in Figure 14. Towards ultra-low bitrates like 0.011 bpp where the information is severely distorted, we find increasing the learning rate of fake score network training can better capture the distorted distribution to enhance realism, as shown in Table 1. Several visual examples are presented in Figure 15, where the CoD-based DMD loss markedly reduces artifacts and yields clearer edges and finer details. Since the encoder is fixed, the encoded information remains heavily distorted under MSE-only optimization, resulting in poor overall reconstruction. We expect that jointly tuning the entire network with the CoD-based DMD loss could further improve perceptual quality.
@@
Figure 17: Evaluating $\mathcal{V}$-and $\mathcal{X}$-prediction pixel-space CoD using DiffC on Kodak at $512\times 512$.
Appendix D Analysis on Different Downstream Codecs
In this section, we further demonstrate the capability of CoD as a foundation model through comparing different downstream codecs in Figure 16.
D.1 Comparison of Downstream Codecs
Zero-shot DiffC. No perceptual optimization is involved. It yields the highest PSNR and strong perceptual quality, but has slow encoding speed. Runtime grows with bitrate, and encoding a 4 bpp image can take up to 100 seconds.
Finetuned one-step CoD. This scheme attains the best perceptual scores, especially DISTS and FID. The single-step diffusion process enables very fast encoding. However, end-to-end perceptual finetuning with perceptual losses leads to relatively lower PSNR.
CoD-based perceptual optimization for MS-ILLM. Although slightly worse than zero-shot DiffC, it requires a lightweight model and offers substantially faster coding, providing a different balance between coding speed and compression ratio.
D.2 Finetuning Multi-step CoD at Higher Bitrates
A straightforward downstream compression scheme is to directly finetune multi-step CoD at higher bitrates through diffusion loss. However, this leads to inferior results compared to other schemes. When jointly optimizing a condition under explicit diffusion loss, the optimization becomes biased: the condition tends to maximize the score norm, as shown in [max_score_norm]. In practice, this drives the decoder toward oversharpened and over-saturated reconstructions, resulting in lower fidelity. This observation is consistent with recent findings that one-step diffusion codecs perform better because they avoid this biased optimization.
The tendency toward a higher score norm also motivates training the foundation model at ultra-low bitrates. Under a strong information bottleneck, the condition can only preserve a small amount of information. The distortion is large at these rates, forcing the available condition to focus on reducing diffusion loss rather than amplifying the score norm. At higher bitrates, reducing diffusion loss becomes easier, leaving room for the condition to optimize for score norm instead. In addition, our unified training strategy and auxiliary losses further help counteract this bias by providing additional compression supervision to the condition network.
Appendix E Exploring $\mathcal{X}$-Prediction Pixel Diffusion
@@
Method @ 0.0039 bpp PSNR LPIPS DISTS FID
CoD (pixel) 16.21 dB 0.434 0.186 46.0
CoD (pixel-$\mathcal{X}$) 17.29 dB 0.416 0.179 44.8
CoD (pixel-$\mathcal{X}$, JiT) 16.89 dB 0.422 0.183 47.4
Table 2: Ablation study for $\mathcal{X}$-prediction on Kodak at $512\times 512$.
Comparison on pixel-space CoD. We follow the JiT training pipeline to build a pixel-space CoD variant using $\mathcal{X}$-prediction and $\mathcal{V}$-loss, denoted as CoD (pixel-$\mathcal{X}$) to distinguish it from our velocity-based pixel-space CoD. To further examine the influence of network design (i.e., the decoupled head [ddt, pixnerd]) we also substitute our pixel-space diffusion model with JiT’s DiT backbone, referred to as CoD (pixel-$\mathcal{X}$, JiT). As shown in Table 2, $\mathcal{X}$-prediction yields significantly better perceptual metrics than $\mathcal{V}$-prediction. In contrast, replacing the decoupled-head design with JiT’s pure DiT structure does not provide performance gains.
Evaluation on downstream DiffC. In Figure 17, we evaluate CoD (pixel-$\mathcal{X}$) on the downstream codec DiffC. Although $\mathcal{X}$-prediction improves performance on the foundation model, it does not consistently benefit DiffC. This is because converting the predicted $x$ into velocity $v$ at time $t$ requires clamping the denominator to be no smaller than 0.05 for stability (following JiT): $v=(x-x_{t})\ /\ \textrm{clamp}(1-t,0.05,1)$. This clamping leads to inaccurate likelihood estimation near $t=1$. When applied to DiffC, the resulting reconstructions retain slight noisy, i.e., the inversion process does not fully denoise the image, leading to degraded FID. The effect is more pronounced at very high bitrates, where the inaccurate likelihood directly impacts DiffC and causes all metrics to drop relative to $\mathcal{V}$-prediction. These results highlight that different prediction targets offer different strengths, and the choice should be adapted to the task objective.
Appendix F Additional Details and Results